# How Safe Will I Be Given What I Saw? Calibrated Prediction of Safety Chances for Image-Controlled Autonomy

### v2

@@ 2.3 World Models
For systems with high-resolution images and complex dynamical models, trajectory prediction becomes particularly challenging. Physics-based methods and classical machine learning methods may not be applicable to these tasks, leading to the almost exclusive application of deep learning for prediction. Sequence prediction models, which originate in deep video prediction Oprea et al. (2022), often struggle in our setting due to compounding prediction errors, the difficulty of modeling long-term dependencies, and the large data requirements for learning high-dimensional dynamics.. However, incorporating additional information, such as states, observations, and actions, in sequence predictions can improve their performance Oh et al. (2015); Finn and Levine (2017); Strickland et al. (2018) — and our predictors take advantage of that. Another way to improve performance is to reduce the input dimensionality. High-dimensional sensor data often contains substantial redundant and irrelevant information, which leads to higher computational costs. Generative adversarial networks (GANs) Goodfellow et al. (2020) map the observations into a low-dimensional latent space, which can enable model-based assurance algorithms Katz et al. (2021), and low-dimensional representations can incorporate conformal predictions to monitor the real-time performance Boursinos and Koutsoukos (2021).
World models Ha and Schmidhuber (2018) are a model-based approach in which a learned environment model serves as a surrogate for the real environment, allowing the controller to be trained largely in imagination rather than through direct interaction with the physical world. This can substantially improve sample efficiency and stability compared to training solely in the real environment. They achieve better results in training the controllers than basic RL. The world model learns an approximation of the system’s dynamics by collecting observations and actions as inputs and outputs the future states. The observation encoder, typically implemented as a VAE, compresses each observation into a latent space (low-dimensional vectors) and trains a mixed-density recurrent neural network (MDN-RNN) to forecast the next-step latent observations. The decoder of the VAE maps the latent vector into the observation space. This VAE–MDN-RNN world model is then applied to do trajectory predictions and get the competency assessment through the forecasting trajectory Acharya et al. (2022).
@@
Recent advances explore the use of discrete latent spaces for more structured and sample-efficient planning, including transformer-based discrete world models for continuous control Scannell et al. (2025), building on neural discrete representation learning Van Den Oord et al. (2017) and echoing our use of VQVAE-based encoders. Building on the original VAE–MDN-RNN architecture of world models Ha and Schmidhuber (2018), more advanced models like DreamerV2 Hafner et al. (2022) and IRIS Micheli et al. (2023) adopt recent deep computer vision models like vision transformers to achieve more vivid images. Transformers have also been shown to improve sample efficiency in latent model training, a crucial factor for model-based planning and control Micheli et al. (2022). Recent systems such as DayDreamer Wu et al. (2023) extend this line of work by leveraging compact world models for sample-efficient control in physical robots, demonstrating the practicality of Dreamer-style pipelines. More recently, Orbis Mousakhan et al. (2025) compares different latent spaces within a unified hybrid tokenizer, showing that a flow-matching-based continuous model achieves superior long-horizon prediction in challenging driving scenarios using only raw video data.
@@
The trajectories predicted by world models have been used to provide uncertainty measurements for agents Acharya et al. (2023), albeit without statistical guarantees so far. In our method, we adopt different architectures of world models and compare them with other one-step monolithic deep learning models using the recurrent modules. Building on this foundation, recent work has demonstrated that pre-trained, interpretable world models can enable zero-shot safety prediction from visual inputs Mao et al. (2024a), and that weakly supervised latent representations can further improve physical consistency and trajectory-level accuracy Mao and Ruchkin (2024). These efforts are aligned with emerging principles for physically interpretable world models Peper et al. (2025), which call for organized latent spaces, aligned invariant/equivariant representations, integration of diverse supervision, and partitioned outputs for scalability. Compared with these approaches, this article develops chance calibration guarantees, which were previously unexplored.
@@ 2.4 Unsupervised Domain Adaptation
Predicting future observations from high-dimensional sensory inputs, such as image sequences, inevitably leads to distribution shift, especially over long prediction horizons. This shift may arise from subtle environmental variations, motion blur, sensor noise, or compounding prediction errors. When applied to such shifted predictions, safety evaluators often produce unreliable or overconfident outputs, which compromise downstream decision-making Yang et al. (2024); Stocco et al. (2020). To address this, we adopt a principled solution of unsupervised domain adaptation (UDA) that continuously adapts the safety evaluator to the evolving input distribution without requiring manual relabeling. This enables the system to maintain calibration guarantees even as the environment changes or predictions drift from training conditions.
A range of UDA strategies has been developed across machine learning and robotics. Early methods, such as Deep CORAL Sun and Saenko (2016) and MMD-based alignment Long et al. (2015), align source and target feature distributions via statistical matching. Adversarial approaches like DANN Ganin et al. (2016) and ADDA Tzeng et al. (2017) train domain-invariant representations by using domain discriminators. Others operate at the image level using GAN-based transformations Bousmalis et al. (2017); Zhu et al. (2017); Sobolewski et al. (2025). More recent work enables adaptation at test time, including entropy-based methods like Tent Wang et al. (2020) and contrastive self-supervised approaches such as CoSCA Kang et al. (2019). In our work, we adopt a lightweight method using the entropy minimization (MEMO) approach Zhang et al. (2022), which does not require source data during adaptation. It is well-suited for safety evaluation in long-horizon prediction settings because it can be applied efficiently within the prediction pipeline without querying external data.
@@ 3.1 Problem Setting
Figure 2: Comparison of safety prediction pipelines and their associated world model structures (gray boxes). Top: The monolithic predictor maps image sequences directly to safety probabilities $P(\varphi(x))$, followed by a calibrator for improved confidence. Middle: The latent monolithic predictor encodes images into latent states and predicts safety from these, with calibration applied afterward. Bottom: The composite latent predictor encodes inputs, forecasts future latent states, and uses an evaluator with UDA to estimate safety. All pipelines predict the probability of satisfying the safety property $\varphi(x)$.
@@
We focus on robotic systems with observation spaces with thousands of pixels and unknown non-linear dynamical and observation models. In such systems, while the state space $\mathbf{X}$ is conceptually known (if only to define $\varphi$), it is not necessary (and often difficult) to construct $f$ and $o$ because the controller acts directly on the observation space $\mathbf{Y}$. Without relying on $f$ and $o$, end-to-end methods like deep reinforcement learning Mnih et al. (2015) and imitation learning Hussein et al. (2017) are used to train a controller h by using data from the observation space. Once the controller is deployed in state $x_{0}$, the system executes a trajectory, which is a sequence $\{x_{i},y_{i},u_{i}\}^{t}_{i=0}$ up to time $t$, where:
@@
Our goal is to predict the system’s safety $\varphi(x_{i+k})$ at time $i+k$ given a series of $m$ observations $\mathbf{y}_{i}=(y_{i-m+1},...,y_{i})$. This sequence of observations does not, generally, determine the true state $x_{i}$ (e.g., when $m=1$, the function $o$ may not be invertible). This leads to partial state observability, which we model stochastically. Specifically, we say that from the predictor’s perspective, $x_{i}$ is drawn from some belief distribution $\mathcal{D}_{\mathbf{y}_{i}}$. This induces a distribution of subsequent trajectories and transforms future safety $\varphi(x_{i+k})$ into a Bernoulli random variable. Therefore, we will estimate the conditional probability $P(\varphi(x_{i+k})\mid\mathbf{y}_{i})$ and provide an error bound on our estimates. Note that process noise in $f_{d}$ and measurement noise in $f_{o}$ are orthogonal to the issue of partial observability; nonetheless, both of these noises are supported by our approach and would be treated as part of the stochastic uncertainty in the future $\varphi$. To sum up the above, we arrive at the following problem description.
@@ 3.2 Predictors and Datasets
Both types of predictors are trained using labeled observation-action datasets, which combine each observation with the controller’s executed action and the eventual safety label. This setup enables training without sacrificing predictive fidelity, as actions encode critical behavioral information even in visually ambiguous settings. All models described in this article are trained on datasets of this form.
@@ 4.1 Monolithic and Composite Predictors
Composite pipelines explicitly separate forecasting from evaluation. In the image space, a forecaster $P_{g}$ predicts future frames that are then classified by a CNN. In the latent space, a forecaster $P_{l}$ outputs $(\theta^{\prime}_{i+1},\ldots,\theta^{\prime}_{i+k})$, after which $V$ inspects the final (or entire) latent trajectory. Similarly, nothing in the design of the composite pipeline restricts how the forecaster is built; it may rely on recurrence, attention, or any hybrid thereof.
@@
For the latent forecaster $P_{l}$, we compare three architectures. The first is a unidirectional LSTM, which provides a strong recurrent baseline. The second extends this with a bidirectional LSTM that incorporates a backward pass, allowing predictions during training to condition on both preceding and succeeding elements of the input sequence Schuster and Paliwal (1997). This enables the model to construct richer temporal representations with enhanced coherence and accuracy of the forecasted output window. Lastly, we adopt a Transformer-based architecture in which each layer combines multi-head self-attention with a position-wise feed-forward network, connected via residual links and layer normalization Vaswani et al. (2017). This design allows every time step in the input window to attend to all others within the same window, enabling the model to capture complex long-range dependencies more effectively than recurrent models. Once the forecast $(\hat{\theta}_{i+k-n+1},\dots,\hat{\theta}_{i+k})$ is produced, the evaluator $V$ emits only a single binary label and does not benefit from future context; therefore, we implement $V$ as a lightweight unidirectional LSTM, which is both parameter-efficient and sufficient for this one-step task.
In the composite latent predictor pipeline, in addition to the standard continuous-latent VAE encoder, we also evaluate a discrete-latent variant based on the Vector Quantized VAE (VQ-VAE) architecture Van Den Oord et al. (2017). The VQ-VAE maps each observation into a sequence of discrete codes drawn from a learned codebook, which serves as a compact, structured representation of the input. This comparison allows us to assess whether discrete latent representations offer improved stability and safety-prediction accuracy in the composite pipeline.
@@ 4.2 Training Process
0:Input: A validation dataset bin $B=\{b_{k}\}_{k=1,\dots}$, each containing sequences of observations and safety $b_{k}=(y_{k},\phi(x_{k}))$, trained safety chance predictor $g$, and miscoverage level $\alpha$
0:Output: confidence bound $c$ satisfying Equation. 4.
@@
2: for $j=1$ to $Q$ do
3:  for $i=1$ to $M$ do
4:   $B_{i}\leftarrow N$ i.i.d. samples from $B$ {resampled bin}
5:   $q_{i}\leftarrow\frac{1}{N}\sum_{l=1}^{N}g(y_{l}),\;\text{for each }y_{l}\in B_{i}$ {mean safety chance prediction}
6:   $p_{i}\leftarrow\frac{1}{N}\sum_{l=1}^{N}\phi(x_{l}),\;\text{for each }\phi(x_{l})\in B_{i}$ {true safety chance}
7:   $\delta_{i}\leftarrow\lvert q_{i}-p_{i}\rvert$ {non‐conformity score}
8:  end for
9:  $n\leftarrow\lceil(M+1)(1-\alpha)\rceil$ {conformal quantile}
10:  $c\leftarrow$ the nn‐th smallest value among $\{\delta_{1},\ldots,\delta_{M}\}$
11: end for
12: return $c$
@@ 4.3 Unsupervised Domain Adaptation
We adopt the Marginal Entropy Minimization (MEMO) approach Zhang et al. (2022) for per-sample test-time adaptation. The key idea is to encourage the evaluator to produce confident and consistent predictions across multiple stochastically augmented versions of the same observation. This is achieved by: (1) generating augmented versions of the input visual observations; (2) aggregating their predictions into a marginal probability distribution; and (3) minimizing the entropy of this marginal distribution by updating the evaluator parameters.
@@
As indicated in Figures 1 and  2, this label-free, per-sample adaptation procedure improves evaluator robustness under latent distribution shift in composite latent predictors.
@@ 4.4 Conformal Calibration for Chance Predictors
To turn a label predictor $\rho$ into a chance predictor $g$, we take its normalized softmax outputs and perform post‐hoc calibration  Guo et al. (2017); Zhang et al. (2020). On a held‐out calibration dataset $Z_{cal}$, we hyperparameter‐tune over state‐of‐the‐art post‐hoc calibration techniques: temperature scaling, logistic calibration, beta calibration, histogram binning, isotonic regression, ensemble of near‐isotonic regression (ENIR), and Bayesian binning into quantiles (BBQ) Wang (2023).
To evaluate the quality of probability estimates produced by our chance predictors, we consider two standard calibration metrics: Expected Calibration Error (ECE) (Guo et al., 2017) and the Brier Score (Glenn et al., 1950). Let $\hat{p}_{i}\in[0,1]$ denote the predicted probability of safety for sample $i$, and $y_{i}\in\{0,1\}$ the corresponding ground truth label. Given $n$ samples and a binning function $\mathrm{bin}(\cdot)$ that assigns predictions to $Q$ disjoint bins, ECE is defined as
@@
To provide this guarantee, we apply conformal prediction (Lei et al., 2018) in Algorithm 1. Intuitively, we rank the safety chance errors in resampled bins and obtain a statistical upper bound on this error for each bin. Similarly to existing works relying on conformal prediction (Lindemann et al., 2023; Fan et al., 2020), this algorithm guarantees (9).
Theorem 1 (Conformal Bounds on Confidence Scores (adaptation of Theorem 2.1 in (Lei et al., 2018))).
@@
The proof follows the same structure as Theorem 2.1 in (Lei et al., 2018), with the modification that non-conformity scores are computed on average safety probabilities within bins. Given that resampled bins $B_{j}$ are drawn i.i.d. from $\mathcal{D}$, the exchangeability assumption required by conformal prediction still holds. Therefore, ranking the absolute deviations $\delta_{j}$ yields the $(1-\alpha)$ coverage guarantee in (9). ∎
Table 1: F1 scores of evaluators on different datasets (in %).
Case Study Evaluator Type Normal OOD Mix
Racing Car Image Evaluator 99.56 63.94 78.23
Latent Evaluator 98.79 65.73 80.27
UDA Image Evaluator 97.13 81.65 91.34
Cart Pole Image Evaluator 99.66 67.21 91.05
Latent Evaluator 99.67 69.14 89.33
UDA Image Evaluator 99.06 76.52 96.62
Donkey Car Image Evaluator 96.20 93.03 94.43
Latent Evaluator 97.19 92.21 94.70
UDA Image Evaluator 96.05 94.14 95.97
Table 2: FPR of evaluators on different datasets (in %).
Case Study Evaluator Type Normal OOD Mix
Racing Car Image Evaluator 0.42 52.78 23.53
Latent Evaluator 2.15 30.82 19.62
UDA Image Evaluator 0.62 7.07 4.37
Cart Pole Image Evaluator 0.53 21.51 9.88
Latent Evaluator 0.61 18.34 7.69
UDA Image Evaluator 1.34 16.72 10.03
Donkey Car Image Evaluator 2.69 18.25 6.05
Latent Evaluator 3.05 11.97 5.32
UDA Image Evaluator 3.21 9.15 5.42

### v3

@@ 2.3 World Models
For systems with high-resolution images and complex dynamical models, trajectory prediction becomes particularly challenging. Physics-based methods and classical machine learning methods may not be applicable to these tasks, leading to the almost exclusive application of deep learning for prediction. Sequence prediction models, which originate in deep video prediction, often struggle in our setting due to compounding prediction errors, the difficulty of modeling long-term dependencies, and the large data requirements for learning high-dimensional dynamics [Oprea_2022]. However, incorporating additional information, such as states, observations, and actions, in sequence predictions can improve their performance — and our predictors take advantage of that [oh2015action, finn2017deep, strickland2018deep]. Another way to improve performance is to reduce the input dimensionality. High-dimensional sensor data often contains substantial redundant and irrelevant information, which leads to higher computational costs. Generative adversarial networks (GANs) map the observations into a low-dimensional latent space, which can enable model-based assurance algorithms, and low-dimensional representations can incorporate conformal predictions to monitor the real-time performance [goodfellow2020generative, katz2021verification, boursinos_assurance_2021]. World models [ha_recurrent_2018] are a model-based approach in which a learned environment model serves as a surrogate for the real environment, allowing the controller to be trained largely in imagination rather than through direct interaction with the physical world. This can substantially improve sample efficiency and stability compared to training solely in the real environment. They achieve better results in training the controllers than basic RL. The world model learns an approximation of the system’s dynamics by collecting observations and actions as inputs and outputs the future states. The observation encoder, typically implemented as a VAE, compresses each observation into a latent space (low-dimensional vectors) and trains a mixed-density recurrent neural network (MDN-RNN) to forecast the next-step latent observations. The decoder of the VAE maps the latent vector into the observation space. This VAE–MDN-RNN world model is then applied to do trajectory predictions and get the competency assessment through the forecasting trajectory [acharya_competency_2022].
@@
Recent advances explore the use of discrete latent spaces for more structured and sample-efficient planning, including transformer-based discrete world models for continuous control, building on neural discrete representation learning and echoing our use of VQVAE-based encoders [scannell2025discrete, van2017neural]. Building on the original VAE–MDN-RNN architecture of world models, more advanced models like DreamerV2 and IRIS adopt recent deep computer vision models like vision transformers to achieve more vivid images [ha_recurrent_2018, hafner2022mastering, micheli2023transformers]. Transformers have also been shown to improve sample efficiency in latent model training, a crucial factor for model-based planning and control [micheli2022transformers]. Recent systems such as DayDreamer extend this line of work by leveraging compact world models for sample-efficient control in physical robots, demonstrating the practicality of Dreamer-style pipelines [wu2023daydreamer]. More recently, Orbis compares different latent spaces within a unified hybrid tokenizer, showing that a flow-matching-based continuous model achieves superior long-horizon prediction in challenging driving scenarios using only raw video data [mousakhan2025orbis].
@@
The trajectories predicted by world models have been used to provide uncertainty measurements for agents, albeit without statistical guarantees so far [acharya2023learning]. In our method, we adopt different architectures of world models and compare them with other one-step monolithic deep learning models using the recurrent modules. Building on this foundation, recent work has demonstrated that pre-trained, interpretable world models can enable zero-shot safety prediction from visual inputs, and that weakly supervised latent representations can further improve physical consistency and trajectory-level accuracy [mao2024zero, mao2024towards]. These efforts are aligned with emerging principles for physically interpretable world models, which call for organized latent spaces, aligned invariant/equivariant representations, integration of diverse supervision, and partitioned outputs for scalability [peper2025four]. Compared with these approaches, this article develops chance calibration guarantees, which were previously unexplored.
@@ 2.4 Unsupervised Domain Adaptation
Predicting future observations from high-dimensional sensory inputs, such as image sequences, inevitably leads to distribution shift, especially over long prediction horizons. This shift may arise from subtle environmental variations, motion blur, sensor noise, or compounding prediction errors. When applied to such shifted predictions, safety evaluators often produce unreliable or overconfident outputs, which compromise downstream decision-making [yang2024generalized, stocco2020misbehaviour]. To address this, we adopt a principled solution of unsupervised domain adaptation (UDA) that continuously adapts the safety evaluator to the evolving input distribution without requiring manual relabeling. This enables the system to maintain calibration guarantees even as the environment changes or predictions drift from training conditions.
A range of UDA strategies has been developed across machine learning and robotics. Early methods, such as Deep CORAL and MMD-based alignment, align source and target feature distributions via statistical matching [sun2016deep, long2015learning]. Adversarial approaches like DANN and ADDA train domain-invariant representations by using domain discriminators [ganin2016domain, tzeng2017adversarial]. Others operate at the image level using GAN-based transformations [bousmalis2017unsupervised, zhu2017unpaired, sobolewski2025generalizable]. More recent work enables adaptation at test time, including entropy-based methods like Tent and contrastive self-supervised approaches such as CoSCA [wang2020tent, kang2019contrastive]. In our work, we adopt a lightweight method using the entropy minimization (MEMO) approach, which does not require source data during adaptation [zhang2022memo]. It is well-suited for safety evaluation in long-horizon prediction settings because it can be applied efficiently within the prediction pipeline without querying external data.
@@ 3.1 Problem Setting
We focus on robotic systems with observation spaces with thousands of pixels and unknown non-linear dynamical and observation models. In such systems, while the state space $\mathbf{X}$ is conceptually known (if only to define $\varphi$), it is not necessary (and often difficult) to construct $f$ and $o$ because the controller acts directly on the observation space $\mathbf{Y}$. Without relying on $f$ and $o$, end-to-end methods like deep reinforcement learning and imitation learning are used to train a controller h by using data from the observation space  [mnih2015human, hussein_imitation_2017]. Once the controller is deployed in state $x_{0}$, the system executes a trajectory, which is a sequence $\{x_{i},y_{i},u_{i}\}^{t}_{i=0}$ up to time $t$, where:
@@
Our goal is to predict the system’s safety $\varphi(x_{i+k})$ at time $i+k$ given a series of $m$ observations $\mathbf{y}_{i}=(y_{i-m+1},...,y_{i})$. This sequence of observations does not, generally, determine the true state $x_{i}$ (e.g., when $m=1$, the function $o$ may not be invertible). This leads to partial state observability, which we model stochastically. Specifically, we say that from the predictor’s perspective, $x_{i}$ is drawn from some belief distribution $\mathcal{D}_{\mathbf{y}_{i}}$. This induces a distribution of subsequent trajectories and transforms future safety $\varphi(x_{i+k})$ into a Bernoulli random variable. Therefore, we will estimate the conditional probability $P(\varphi(x_{i+k})\mid\mathbf{y}_{i})$ and provide an error bound on our estimates. Note that process noise in $d$ and measurement noise in $o$ are both supported by our approach and would be treated as part of the stochastic uncertainty in the future $\varphi$. To sum up the above, we arrive at the following problem description.
@@ 3.2 Predictors and Datasets
Both types of predictors are trained using labeled observation-action datasets, which combine each observation with the controller’s executed action and the eventual safety label. This setup enables training without sacrificing predictive fidelity, as actions encode critical behavioral information even in visually ambiguous settings. Thus, all models described in this article are trained on datasets of the following form:
@@
Notice that our dataset is structured not around sequential states within a single trajectory, but as independent trajectories sampled from the same distribution. This ensures that the calibration and test samples are i.i.d., satisfying the exchangeability requirement of conformal prediction. We are not applying conformal prediction to temporally correlated states along a single trajectory; rather, each data point represents an independent realization of the same prediction task.
@@ 4.1 Monolithic and Composite Predictors
Figure 2: Comparison of safety prediction pipelines and their world-model structures (gray boxes). Top: The monolithic predictor maps image sequences directly to safety probabilities $P(\varphi(x))$, followed by a calibrator for improved confidence. Middle: The latent monolithic predictor encodes images into latent states and predicts safety from these, with calibration applied afterward. Bottom: The composite latent predictor encodes inputs, forecasts future latent states, and uses an evaluator with UDA to estimate safety. All pipelines predict the probability of satisfying the safety property $\varphi(x)$.
@@
Composite pipelines explicitly separate forecasting from evaluation. In the image space, a forecaster $P_{g}$ predicts future frames that are then classified by a CNN. In the latent space, a forecaster $P_{l}$ outputs $(\theta^{\prime}_{i+1},\ldots,\theta^{\prime}_{i+k})$, after which $V$ inspects the latent trajectory. Similarly, the design of the composite pipeline does not restrict how the forecaster is built; it may rely on recurrence, attention, or any hybrid thereof.
@@
For the latent forecaster $P_{l}$, we compare three architectures. The first is a unidirectional LSTM, which provides a strong recurrent baseline. The second extends this with a bidirectional LSTM that incorporates a backward pass, allowing predictions during training to condition on both preceding and succeeding elements of the input sequence [schuster_bidirectional_1997]. This enables the model to construct richer temporal representations with enhanced coherence and accuracy of the forecasted output window. Lastly, we adopt a Transformer-based architecture in which each layer combines multi-head self-attention with a position-wise feed-forward network, connected via residual links and layer normalization [vaswani_attention_2017]. This design allows every time step in the input window to attend to all others within the same window, enabling the model to capture complex long-range dependencies more effectively than recurrent models. Once the forecast $(\hat{\theta}_{i+k-n+1},\dots,\hat{\theta}_{i+k})$ is produced, the evaluator $V$ emits only a single binary label and does not benefit from future context; therefore, we implement $V$ as a lightweight unidirectional LSTM, which is both parameter-efficient and sufficient for this one-step task.
In the composite latent predictor pipeline, in addition to the standard continuous-latent VAE encoder, we also evaluate a discrete-latent variant based on the Vector Quantized VAE (VQ-VAE) architecture [van2017neural]. The VQ-VAE maps each observation into a sequence of discrete codes drawn from a learned codebook, which serves as a compact, structured representation of the input. This comparison allows us to assess whether discrete latent representations offer improved stability and safety-prediction accuracy in the composite pipeline.
@@ 4.2 Training Process
1:A validation dataset bin $B=\{b_{k}\}_{k=1,\dots}$, each containing sequences of observations and safety $b_{k}=(y_{k},\phi(x_{k}))$, trained safety chance predictor $g$, and miscoverage level $\alpha$
2:confidence bound $c$ satisfying Equation. 4.
@@
4:for $j=1$ to $Q$ do
5:  for $i=1$ to $M$ do
6:   $B_{i}\leftarrow N$ i.i.d. samples from $B$ {resampled bin}
7:   $q_{i}\leftarrow\frac{1}{N}\sum_{l=1}^{N}g(y_{l}),\;\text{for each }y_{l}\in B_{i}$ {mean safety chance prediction}
8:   $p_{i}\leftarrow\frac{1}{N}\sum_{l=1}^{N}\phi(x_{l}),\;\text{for each }\phi(x_{l})\in B_{i}$ {true safety chance}
9:   $\delta_{i}\leftarrow\lvert q_{i}-p_{i}\rvert$ {non‐conformity score}
10:  end for
11:  $n\leftarrow\lceil(M+1)(1-\alpha)\rceil$ {conformal quantile}
12:  $c\leftarrow$ the nn‐th smallest value among $\{\delta_{1},\ldots,\delta_{M}\}$
13:end for
14:return $c$
@@ 4.3 Unsupervised Domain Adaptation
We adopt the Marginal Entropy Minimization (MEMO) approach for per-sample test-time adaptation [zhang2022memo]. The key idea is to encourage the evaluator to produce confident and consistent predictions across multiple stochastically augmented versions of the same observation. This is achieved by: (1) generating augmented versions of the input visual observations; (2) aggregating their predictions into a marginal probability distribution; and (3) minimizing the entropy of this marginal distribution by updating the evaluator parameters.
@@
As indicated in Figures 1 and 2, this label-free, per-sample adaptation improves evaluator robustness under latent distribution shift in composite latent predictors.
It is important to distinguish between two sources of distribution shift in our setting. First, the decoded images produced by the VAE may themselves be out-of-distribution for the safety evaluator, even when the underlying trajectory is in-distribution. This shift occurs because reconstruction artifacts and latent-space forecasting errors introduce visual distortions absent from the original training images. Second, the trajectories themselves may be out-of-distribution, representing scenarios not encountered during training. Our test-time adaptation strategy addresses the first source directly: by encouraging consistent, confident predictions across stochastically augmented views of each decoded image, MEMO makes the evaluator robust to the visual perturbations introduced by the decode and predict pipeline, without requiring any additional safety labels.
To mitigate the impact of distribution shifts occurring at the trajectory level, our framework currently relies on the conformal calibration mechanism described in the following section. When the system encounters a trajectory that significantly deviates from the training distribution, the predicted safety chance $P^{\prime}(\varphi(x_{i+k})|y_{i})$ naturally reflects higher uncertainty, leading to wider conformal intervals that alert the controller to potential unreliability [yang2024generalized].
This robustness can be further augmented by leveraging pre-trained foundation world models and multimodal meaning representations [mao2024zero, liu2023meaning]. By utilizing the superior zero-shot generalization of these large-scale models, the safety evaluator can potentially reason about novel maneuvers or environmental geometries not present in the local training set, thereby bridging the gap between training and deployment distributions.
@@ 4.4 Conformal Calibration for Chance Predictors
To turn a label predictor $\rho$ into a chance predictor $g$, we take its normalized softmax outputs and perform post‐hoc calibration [guo_calibration_2017, zhang_mix-n-match_2020]. On a held‐out calibration dataset $Z_{cal}$, we hyperparameter‐tune over state‐of‐the‐art post‐hoc calibration techniques: temperature scaling, logistic calibration, beta calibration, histogram binning, isotonic regression, ensemble of near‐isotonic regression (ENIR), and Bayesian binning into quantiles (BBQ) [wang2023calibration].
To evaluate the quality of probability estimates produced by our chance predictors, we consider two standard calibration metrics: Expected Calibration Error (ECE) and the Brier Score [glenn1950verification, guo_calibration_2017]. Let $\hat{p}_{i}\in[0,1]$ denote the predicted probability of safety for sample $i$, and $y_{i}\in\{0,1\}$ the corresponding ground truth label. Given $n$ samples and a binning function $\mathrm{bin}(\cdot)$ that assigns predictions to $Q$ disjoint bins, ECE is defined as
@@
To provide this guarantee, we apply conformal prediction in Algorithm 1 [lei2018distribution]. Intuitively, we rank the safety chance errors in resampled bins and obtain a statistical upper bound on this error for each bin. Similarly to existing works relying on conformal prediction, this algorithm guarantees equation (9[lindemann_conformal_2023, fan2020statistical]. Recall that each calibration example $(\mathbf{y}_{i},\varphi(x_{i+k}))$ is drawn from an independent trajectory, so the calibration set satisfies the i.i.d. assumption required by conformal prediction.
Theorem 1 (Conformal Bounds on Confidence Scores (adaptation of Theorem 2.1 in Lei et al. 2018[lei2018distribution])).
@@
The proof follows the same structure as Theorem 2.1 in Lei et al. 2018, with the modification that non-conformity scores are computed on average safety probabilities within bins[lei2018distribution]. Given that resampled bins $B_{j}$ are drawn i.i.d. from $\mathcal{D}$, the exchangeability assumption required by conformal prediction still holds. Therefore, ranking the absolute deviations $\delta_{j}$ yields the $(1-\alpha)$ coverage guarantee in (9). ∎